# Helmlab: A Data-Driven Analytical Color Space for UI Design Systems

### v1

@@ Abstract
We present Helmlab, a 72-parameter analytical color space for UI design systems. The forward transform maps CIE XYZ to a perceptually-organized Lab representation through learned matrices, per-channel power compression, Fourier hue correction, and embedded Helmholtz–Kohlrausch lightness adjustment. A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries. On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.22, a 20.4% reduction from CIEDE2000 (29.18). Cross-validation on He et al. 2022 and MacAdam 1974 shows competitive cross-dataset performance. The transform is invertible with round-trip errors below $10^{-14}$. Gamut mapping, design-token export, and dark/light mode adaptation utilities are included for use in web and mobile design systems.
@@ 1 Introduction
    A 72-parameter analytical color space with stress 23.22 on combvd, 20.4% lower than CIEDE2000.
@@
    A rigid rotation in the $ab$ plane that reduces hue error (RMS 16.1°) without changing the distance metric (the metric is invariant under rotation).
@@ 3.1 Forward Transform
where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation. The optimized weight is $w_{\mathrm{HK}}=0.37$ and the power is $p_{\mathrm{HK}}=0.85$, making lightness explicitly dependent on chroma. Most existing spaces treat lightness and chroma as independent.
@@ 4 Distance Metric
The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}\approx 10^{-4}$, $s_{C}\approx 0.0045$). The Minkowski exponent $p=0.83$ and chroma weight $w_{C}=1.03$ slightly favor chroma differences over lightness. The compression $c=1.59$ and post-power $q=1.1$ reduce systematic bias for large color differences.
@@ 5 Optimization
$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}$ (16)
using L-BFGS-B [10] with box constraints. We perform 8 random restarts (each 5,000 iterations) and keep the best. The He et al. term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
@@
The optimizer typically converges within 3,000–5,000 function evaluations per restart. The validation gap (train–val stress difference) is +1.11, which shows some overfitting but not severe for a 72-parameter model.
@@ 6.1 Color-Difference Prediction
Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods. Helmlab achieves 23.22, compared to 29.18 for CIEDE2000 (-20.4%). A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$. The intervals do not overlap ($p<10^{-4}$).
@@
The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression. Oklab’s high stress (47.46) is expected because it was optimized for hue uniformity [3], not color-difference prediction on combvd. The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms. With Euclidean distance only, Helmlab gets stress $\approx$30.2, which is close to CIEDE2000 (29.18). This means the space transform accounts for roughly half the improvement and the distance metric for the other half.
@@ 6.2 Cross-Validation
On He et al. 2022 (regularizer, 82 pairs), Helmlab achieves stress 29.0 vs CIEDE2000’s 32.6. On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.2 vs CAM16-UCS’s 18.7 (best). Figure˜4 visualizes these results. Helmlab outperforms CIEDE2000 on He et al. but trails CAM16-UCS on MacAdam. This is consistent with what one would expect from a model trained primarily on combvd: it generalizes reasonably but does not dominate on all evaluation sets.
@@ 6.3 Generation Properties: The Measurement–Generation Tradeoff
The neutral correction (Section˜3, Stage 10) addresses this problem. It subtracts the achromatic error after all enrichment stages, so the optimizer can find its measurement-optimal parameters without constraint. The achromatic axis is then corrected separately. The stress increase from this correction is +0.04+0.04 (23.18 to 23.22), while achromatic chroma drops from 0.34 to $10^{-6}$.
@@
Val gap (train-val) +1.11
He stress 29.03
MacAdam stress 20.23
The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export. The hue RMS of 16.1° means sRGB primaries land within about 20° of their expected positions. This is acceptable for palette generation but not perfect. The rigid rotation (Section˜3, Stage 11) provides this hue improvement without affecting the distance metric because the metric is invariant under rotation. The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
@@ 6.4 Ablation Study
Table˜3 isolates the contribution of key components.
Table 3: Ablation: effect of removing components.
@@
Full Helmlab 23.22
Euclidean distance only $\sim$30.2 +7.0
No H-K embedding ($w_{\mathrm{HK}}=0$) $\sim$24.5 +1.3
No hue correction $\sim$25.8 +2.6
No dark L compression $\sim$23.9 +0.7
Rotation (-28.2°) 23.22 0.00
No rotation 23.22 0.00
With Euclidean distance, the space alone gets stress 30.2, which is slightly worse than CIEDE2000 (29.18). The full metric brings it down to 23.22, recovering the remaining 7 points. Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone already gives a clear improvement. Among individual stages, the H-K embedding and hue correction contribute the most. The rotation row confirms that the metric is exactly invariant under rotation.
@@ 8 Limitations and Future Work
    Training data. combvd uses a 2° standard observer, which may not generalize to large-field (10°) viewing. The MacAdam cross-validation partially addresses this.
@@
    Hue alignment. RMS hue error of 16.1° is a compromise. Some hue regions (cyan, yellow) have errors up to 20°. Further harmonic terms could improve this at the risk of overfitting.
@@
    Sub-dataset variation. Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (BFD-P(C), Leeds, RIT-DuPont), which may indicate that these subsets benefit more from CIE Lab-based corrections.
@@
For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.
@@ 9 Conclusion
We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.22 on combvd, which is 20.4% lower than CIEDE2000. The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 16.1° RMS without affecting the distance metric. The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis at a cost of only +0.04+0.04 stress points. Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab. An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
@@ Appendix A Parameter Table
Table˜4 lists all 72 optimized parameters grouped by pipeline stage.
Table 4: Optimized Helmlab parameters (v19-NC, $\varphi=-28.2$°).
@@
$\mathbf{M}_{1}$ (9) Row 0 [0.724, 0.238, -0.164] XYZ $\to$ cone-like
Row 1 [-0.325, 1.238, -0.024]
Row 2 [0.101, 0.341, 0.849]
$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.395, 0.421, 0.434 Power compression
$\mathbf{M}_{2}$ (9) Row 0 [-0.425, 0.489, 0.684] Compressed $\to$ Lab
Row 1 [1.931, -2.728, 0.765]
Row 2 [0.009, 1.642, -1.250]
@@
H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.370, 0.852, -0.487, … Chroma $\to$ L
Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.225, -0.748, 0.699, … Cubic + Lh + dark
@@
Distance (7) $s_{L}$ $9.16\times 10^{-5}$ Pair-dep L weight
$s_{C}$ 0.0045 Pair-dep C weight
pp 0.829 Minkowski exponent
$w_{C}$ 1.028 Chroma weight
$c$ 1.588 Compression
@@
Full parameter values are available in the accompanying analytical_params.json file.

### v2

# arXiv:2602.23010v2[cs.GR] 01 Mar 2026

@@ Abstract
We present Helmlab, a 72-parameter analytical color space for UI design systems. The forward transform maps CIE XYZ to a perceptually-organized Lab representation through learned matrices, per-channel power compression, Fourier hue correction, and embedded Helmholtz–Kohlrausch lightness adjustment. A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries. On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.30, a 20.2% reduction from CIEDE2000 (29.18). A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress. Cross-validation on He et al. 2022 and MacAdam 1974 shows competitive cross-dataset performance. The transform is invertible with round-trip errors below $10^{-14}$. Gamut mapping, design-token export, and dark/light mode adaptation utilities are included for use in web and mobile design systems.
@@ 1 Introduction
    A 72-parameter analytical color space with stress 23.30 on combvd, 20.2% lower than CIEDE2000.
@@
    A rigid rotation in the $ab$ plane that reduces hue error (RMS 18.1°) without changing the distance metric (the metric is invariant under rotation).
@@
    A blue-band refit that reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ while improving stress on 5 of 6 combvd sub-datasets.
  6. 6.
@@ 3.1 Forward Transform
where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation. The optimized weight is $w_{\mathrm{HK}}=0.389$ and the power is $p_{\mathrm{HK}}=0.849$, making lightness explicitly dependent on chroma. Most existing spaces treat lightness and chroma as independent.
@@ 4 Distance Metric
The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}=0.001$, $s_{C}=0.022$). The Minkowski exponent $p=0.804$ and chroma weight $w_{C}=1.046$ slightly favor chroma differences over lightness. The compression $c=1.590$ and post-power $q=1.1$ reduce systematic bias for large color differences.
@@ 5 Optimization
$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}+\mathcal{L}_{\mathrm{blue}}$ (16)
using L-BFGS-B [10] with box constraints. We perform 10 random restarts (each 5,000 iterations) and keep the best. The He et al. term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
The blue-band penalty is:
$\mathcal{L}_{\mathrm{blue}}=\lambda_{b}(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+0.5\,\mathrm{STRESS}_{\mathrm{FullBlue}})$ (17)
where $\lambda_{b}=0.1$ and the subscripts denote combvd sub-datasets with high blue-band content. Additional constraints enforce round-trip accuracy ($\lambda_{\mathrm{RT}}=100$), achromatic integrity ($\lambda_{\mathrm{ach}}=500$), and Munsell uniformity ($0.1\cdot\max(\mathrm{CV}-20,0)$). The blue-band penalty was introduced in v2 of this paper to address gradient non-uniformity in the blue–cyan region (Section˜6.6).
@@
The optimizer typically converges within 3,000–5,000 function evaluations per restart. Five-fold cross-validation (with re-optimization per fold) yields a mean validation gap of +0.98+0.98 stress points (std. $0.78$), indicating mild but acceptable overfitting for a 72-parameter model.
@@ 6.1 Color-Difference Prediction
Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods. Helmlab achieves 23.30, compared to 29.18 for CIEDE2000 (-20.2%). A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$. The intervals do not overlap ($p<10^{-4}$).
@@
The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression. Oklab’s high stress (47.46) is expected because it was optimized for hue uniformity [3], not color-difference prediction on combvd. The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms. With Euclidean distance only, Helmlab gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%. The full distance metric then reduces stress by a further 4.3 points. Relative to CIEDE2000, the space transform and the distance metric contribute roughly equally to the total improvement.
@@ 6.2 Cross-Validation
On He et al. 2022 (regularizer, 82 pairs), Helmlab achieves stress 30.3 vs CIEDE2000’s 32.6. On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.3 vs CAM16-UCS’s 18.7 (best). Figure˜4 visualizes these results. Helmlab outperforms CIEDE2000 on He et al. but trails CAM16-UCS on MacAdam. This is consistent with what one would expect from a model trained primarily on combvd: it generalizes reasonably but does not dominate on all evaluation sets.
@@ 6.3 Generation Properties: The Measurement–Generation Tradeoff
The neutral correction (Section˜3, Stage 10) addresses this problem. It subtracts the achromatic error after all enrichment stages, so the optimizer can find its measurement-optimal parameters without constraint. The achromatic axis is then corrected separately. The stress increase from the neutral correction alone is +0.04+0.04 (from a measurement-only baseline of 23.18 to 23.22); the subsequent blue-band refit (Section˜6.6) adds another +0.08+0.08, yielding the final 23.30. Achromatic chroma drops from 0.34 to $10^{-6}$.
@@
5-fold CV gap (mean) +0.98
He stress 30.33
MacAdam stress 20.33
The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export. The hue RMS of 18.1° means sRGB primaries land within about 24° of their expected positions. This is acceptable for palette generation but not perfect. The rigid rotation (Section˜3, Stage 11) provides this hue improvement without affecting the distance metric because the metric is invariant under rotation. The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
@@ 6.4 Ablation Study
Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their optimized values. This measures the marginal contribution of each component given the current parameter setting, not its isolated effect.
Table 3: Ablation: effect of removing components (frozen, not re-optimized).
@@
Full Helmlab 23.30
Euclidean distance only 27.60 +4.3
No H-K embedding ($w_{\mathrm{HK}}=0$) 27.30 +4.0
No hue correction 38.81 +15.5
No dark L compression 23.49 +0.2
Rotation (-28.2°) 23.30 0.00
No rotation 23.30 0.00
With Euclidean distance, the space alone gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%. The full metric brings it down to 23.30, recovering a further 4.3 points. Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone gives a substantial improvement. Among individual stages, hue correction is the most critical ($\Delta=+15.5$), followed by the H-K embedding ($\Delta=+4.0$). These large frozen-ablation deltas reflect the co-optimization of all parameters: disabling one stage disrupts the balance of the entire pipeline. The rotation row confirms that the metric is exactly invariant under rotation.
6.5 Sub-Dataset Performance
Table˜4 breaks down stress by combvd sub-dataset. Helmlab outperforms CIEDE2000 on 3 of 6 sub-datasets (BFD-P(D65), BFD-P(M), Witt). On the blue-heavy sub-datasets (LEEDS, RIT-DuPont), the blue-band refit reduces the gap substantially compared to an earlier version without the penalty: RIT-DuPont improves from 23.49 to 20.41, nearly matching CIEDE2000’s 19.51. BFD-P(C) is the only sub-dataset measured under illuminant C (rather than D65) and with a 2° observer, which may explain the remaining deficit.
Table 4: stress on individual combvd sub-datasets.
Sub-dataset NN Helmlab CIEDE2000
BFD-P(D65) 2,028 23.11 24.14
BFD-P(M) 548 22.22 35.09
BFD-P(C) 200 30.15 30.05
Witt 418 28.88 30.37
Leeds 307 20.92 19.46
RIT-DuPont 312 20.41 19.51
6.6 Blue-Band Analysis
During development, we observed that an initial version of Helmlab (without the blue-band penalty) exhibited severe gradient non-uniformity in the blue–cyan region. A red-to-blue gradient interpolated in Helmlab Lab coordinates showed a $51.4\times$ max/min step-size ratio when measured with CIEDE2000, compared to $1.3\times$ for the same gradient in CIE Lab. Frozen ablation traced the problem to the 4-harmonic hue correction, whose Fourier coefficients overshoot in the blue–magenta band ($h\approx 240$340340°). Disabling hue correction reduced the gradient ratio from $51.4\times$ to $4.9\times$ but increased combvd stress by +15.5+15.5 points, making this an unacceptable trade-off.
The solution was to add explicit sub-dataset penalties to the loss function (Equation˜17) and re-optimize all 72 parameters. With $\lambda_{b}=0.1$, the red-to-blue gradient ratio drops from $51.4\times$ to $5.8\times$ ($8.9\times$ improvement), while combvd stress increases by only +0.08+0.08 (from 23.22 to 23.30). stress improves on 5 of 6 sub-datasets (Table˜4), with the largest gain on RIT-DuPont (3.08-3.08 points). The only regression is BFD-P(D65) (+0.49+0.49), the largest sub-dataset, which accounts for the slight overall stress increase.
@@ 8 Limitations and Future Work
    Training data. combvd predominantly uses the CIE 1964 10° standard observer ($\sim$95% of pairs). Only BFD-P(C) (200 pairs, 5%) uses the 2° observer. Gao et al. 2023 [11] showed that the difference between 2° and 10° observers is not statistically significant for stress evaluation, and MacAdam 1974 (measured with the 2° observer) cross-validation confirms reasonable generalization.
@@
    Hue alignment. RMS hue error of 18.1° is a compromise. Some hue regions (cyan) have errors up to 23.9°. Further harmonic terms could improve this at the risk of overfitting.
@@
    Sub-dataset variation. Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (Table˜4). The blue-band refit improves 5 of 6 sub-datasets compared to the previous version; only BFD-P(C), measured under illuminant C with a 2° observer, remains notably worse (30.15 vs 30.05).
@@
    Blue-band trade-off. The blue-band refit (Section˜6.6) reduces gradient non-uniformity by $8.9\times$ at a cost of +0.08+0.08 stress on combvd. The Munsell CV increases from 17.6% to 20.0%, hue RMS from 16.1° to 18.1°, and the He et al. cross-validation stress increases from 29.0 to 30.3.
  8. 8.
@@
The open-source release includes a second, generation-optimized space (GenSpace) with shared $\gamma\!=\!\tfrac{1}{3}$ and no enrichment stages, which trades distance accuracy for faster, invertible gradient and palette generation. For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.
@@ 9 Conclusion
We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.30 on combvd, which is 20.2% lower than CIEDE2000. The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 18.1° RMS without affecting the distance metric. A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress points, while improving stress on 5 of 6 sub-datasets. The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis. Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab. An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
@@ Appendix A Parameter Table
Table˜5 lists all 72 optimized parameters grouped by pipeline stage.
Table 5: Optimized Helmlab parameters (v20b, $\varphi=-28.2$°).
@@
$\mathbf{M}_{1}$ (9) Row 0 [0.734, 0.240, -0.158] XYZ $\to$ cone-like
Row 1 [-0.330, 1.235, -0.000]
Row 2 [0.086, 0.341, 0.832]
$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.389, 0.416, 0.424 Power compression
$\mathbf{M}_{2}$ (9) Row 0 [-0.420, 0.485, 0.642] Compressed $\to$ Lab
Row 1 [1.933, -2.718, 0.764]
Row 2 [0.006, 1.628, -1.261]
@@
H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.389, 0.849, -0.494, … Chroma $\to$ L
Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.222, -0.727, 0.695, … Cubic + Lh + dark
@@
Distance (7) $s_{L}$ $1.01\times 10^{-3}$ Pair-dep L weight
$s_{C}$ 0.022 Pair-dep C weight
pp 0.804 Minkowski exponent
$w_{C}$ 1.046 Chroma weight
$c$ 1.590 Compression
@@
Full parameter values are available in the accompanying metric_params.json file.